In the blackberry market, the quantity demanded is given by QD = 2,600 – 500P, and the quantity supplied is given by QS = –400 + 100P. What are the equilibrium price and equilibrium quantity?

$5 and 100 pounds

$1.80 and 2,200 pounds

$4.25 and 3,000 pounds

$2.50 and 900 pounds
Answer #1

Setting QD=QS

2600-500P=-400+100P

2600+400=100P+500P

3000=600P

P=3000/600=5

Q=2600-500(5) = 100

Option(A)